# Multitask methods for predicting molecular properties from heterogeneous data

### v1

@@ III Heterogeneous Training Data
In our numerical results, we always train models with the CCSD(T) data from the C set and consider various compositions of the secondary sets. All secondary tasks are highlighted in gray. We identify different training set compositions by the letters of the sets of molecule which are included in our secondary training data. For example, the top diagram in Fig. 2 shows an A training set because the only DFT predictions available to train the model come from the Additional set. By contrast, the lower diagram shows a CAT training set because dots are present in the slots for DFT predictions for Core, Additional, and Target molecular configurations.
Furthermore, the vertical alignment of dots indicate when some of the same molecular configurations are present in the training sets for multiple tasks. In the CAT example in Fig. 2, the dots indicate that the secondary tasks share $20\%$ of the C and A sets while the rest of the molecules in these sets are assigned to only one task each. In Section IV, we will show results for the C, CT, A, AT, CA, and CAT training set configurations. Fig. 6 shows the impact of varying the fraction of molecules that the secondary tasks have in common.
Our testing framework is designed so that models trained with DFT Target data can readily be compared with equivalent models that do not train on this data. Since the inference model can make Target predictions one at a time, for each Target molecule we train a new model using the C set, A set, and the DFT prediction for only that Target molecule. While in practice we could take the much more efficient approach of training on the entire DFT Target set at once, our approach enables us to contrast each CAT model to a CA model with a training set which only differs by one molecule. Consequently, we can isolate the effect of the DFT level Target data on prediction performance.
In Section IV, we will examine the performance of the $\Delta$ and Multitask methods, informed by CCSD(T) and DFT predictions, and compare them to basic Gaussian process regression, informed only by CCSD(T).
@@ IV.1.1 Data Generation
For our first test case—inference of 3-b interaction energy—5986 water trimer configurations and corresponding CCSD(T) calculations were randomly selected from 45,332 configurations available from https://github.com/jmbowma/q-AQUAYu et al. (2022) Density functional theory data was generated for 2992 of these configurations using Psi4. Smith et al. (2020) We employed the PBEPerdew, Burke, and Ernzerhof (1996) and SCANSun, Ruzsinszky, and Perdew (2015) functional approximations. The former was selected for its popularity in materials science research and the latter because its description of intermediate range dispersion has been found to contribute to accurate prediction of water energy differences Chen et al. (2017); Dasgupta et al. (2021) Following recent work on dispersion corrections, Price, Bryenton, and Johnson (2021) we paired the PBE approximation with a D3 correction using finite damping and the SCAN approximation with a D3 correction using zero damping.  Both DFT and CCSD(T) calculations employed a standard counterpoise correction to alleviate basis set superposition error. Huang, Braams, and Bowman (2006)
@@
Quantification of cost in our numerical results refers to the cost of data set generation. The cost of inference is negligible next to quantum chemistry calculations. We base our cost model for each method on the average runtime of Psi4 computations for ten representative systems. For more details about the cost model, see Appendix A.
@@ IV.1.2 Features
We require features which can distinguish between molecular systems to serve as inputs, $\bm{X}_{i}$, to our Gaussian process models. Kernel functions use features to model the relationships between quantities of interest. As noted in Section II.1, in our experiments we consider Smooth Overlap of Atomic Positions (SOAP) features, which have been successfully used to fit interatomic potentials in a Gaussian process based approach.  Bartók et al. (2010); Bartók, Kondor, and Csányi (2013); Bartók and Csányi (2015) SOAP features are constructed based on the neighborhood of each atom in a system, so challenges arise when two systems contain different numbers of constituent atoms of elements. Multiple strategies have been proposed for constructing “global” features to capture entire systems. De et al. (2016) The simplest approach is to take an average over the SOAP features for each atom in the system. Since this approach is computationally efficient and works well in practice, we apply it in our test cases. We chose SOAP parameters following experimentation and the advice of literature.  Deringer et al. (2021); Musil et al. (2021) More details are available in Appendices B and C.
@@ IV.1.3 Hyperparameter Optimization
Squared exponential kernels are used within all Gaussian process priors in our experiments. Abrahamsen (1997) These kernels introduce mean, variance ($v$), and length scale ($\ell$) hyperparameters in addition to the multitask correlation ($\rho$). We use the mean of our training data for each task as that task’s prior mean, calculated at training time. To estimate the remaining parameters, we construct an optimization data set containing $150$ randomly selected molecular configurations which we subsequently exclude from our training and testing data sets. We assume here that we have access to predictions of our quantities of interest from each quantum chemistry method for each of these $150$ molecular configurations, but when this assumption does not hold, a conventional GP regression model may be used to fill in missing data.
We estimate correlation parameters, $\rho$, using Pearson’s correlation coefficient between each secondary task optimization data set and the corresponding primary task data. Popular belief in literature holds that correlation parameters require a detailed optimization procedure, Dral, Zubatiuk, and Xue (2023) but we find that these easy to obtain estimates work reasonably well in practice. A more detailed optimization is left for future work. We estimate $\ell\in\mathbb{R}$ by maximizing its log-marginal likelihood. This approach is easily extended to multifidelity and multitask cases by treating $Y_{ij}-\rho_{j}Y_{i1}$ as the $i^{th}$ observation. Forrester and A. Sóbester (2007) For the organic molecules example, we also use a maximum likelihood estimate for $v$ while for the water trimer example, we find better results from using an estimate of $v$ which minimizes mean absolute prediction error.
@@ IV.2.1 Multitask Performance
Figure 3: Water Trimer Case. Comparison of the MAE of the multitask method for a range of training set constructions (scatter points) to the MAE of a GP regression model trained only with CCSD(T) data (black line). Predictions by the multitask method are informed by CCSD(T) and SCAN. The legends of the plots indicate the structure of the secondary training set according to the system presented by Fig. 2. All MAE reported are the average results obtained from six randomly chosen training sets.
We first compare the performance of multitask models with only one secondary task to the performance of a traditional GP—or single task—model. Fig. 3 shows the error incurred by several implementations of each model when predicting the three-body (3-b) energy of water trimer configurations. We used the SCAN functional to generate the secondary task data, and Pearson’s correlation coefficient between the primary and secondary tasks is $0.997$, indicating strong positive correlation. The three subplots of Fig. 3 cover six secondary training set configurations (C, CT, A, AT, CA, and CAT), labeled within the legends.
All models were tested on a Target set of 320 molecular configurations. We let the size of the Core set, $N_{c}$, take on values in $\{5,10,20,40,80,160,320\}$ for both inference methods and additionally consider $N_{c}\in\{640,1280,2560\}$ for the single task method. The subplots of Fig. 3 also show seven clear groupings of data points with steep slope. This effect is a consequence of the high cost of CCSD(T) relative to DFT. Each of the seven groupings corresponds to a different $N_{c}$. Given the size of the Core set, we let the number of configurations in the A set range from $0.5N_{c}$ to $6N_{c}$. Each point on the plot represents a model with a different training set size. The x axis indicates the cost of the training set while the y axis gives the average mean absolute error for six random assignments of molecular configurations to the C, A, and T sets. We provide details of our cost model in Appendix A.
@@
Figure 4: Water Trimer Case. Effect of including Core and Additional data in secondary training sets. Each indigo point represents a comparison between two multitask models where one model uses a training set which is a subset of the other. In (a), this comparison is between A and CA training configurations, and in (b), we compare AT and CAT configurations. For each pair of models compared, both the primary training data and the secondary data from the A set are identical. The $y=x$ line is plotted in green for reference.
For every training set configuration except C, the multitask method outperforms the single task reference, shown by a black line. Including T data from DFT in the training set leads to better results than corresponding cases without T in every example. In particular, training with secondary T data can enable the multitask method to exhibit an accuracy comparable to a GP model that is an order of magnitude more expensive.
@@
While the C results demonstrate no significant improvement on the single task model, both A and CA results show steady improvement as the size of the A set increases. This behavior suggests that for the multitask method to be beneficial, secondary training data must cover molecular space that is not included in the CCSD(T) training data. We mention that it is possible that the behavior of C models would improve if we consider task dependent features or different relationships between task regression functions. We do not investigate these alternatives further here; instead, we focus on the usual setting where secondary DFT data for additional systems is cheap to obtain.
@@
CT and CAT models are the best-performing multitask models, demonstrating very similar MAE per cost to each other. Furthermore, unlike the AT model, the CAT model does not show improvement as the size of the A set increases. It seems that for our setting, the combination of secondary C and T training data is so valuable to the model that any benefit from A data is dwarfed. We can hypothesize that when the model has access to data for the Core molecules from the primary and secondary tasks, it implicitly learns information about the difference between the tasks. Given secondary data on the Target, it can leverage the captured relationship between tasks to make an accurate prediction for the Target on the primary level.
@@
Fig. 4 provides more insight to the relative roles of the training data sets and to the inherent flexibility of the multitask framework. While Fig. 3 demonstrates that multitask models can be successful when trained on many different training data configurations from the CAT framework, it does not directly compare the performance of the A and CA models. Such a comparison will illuminate the resilience of the multitask model to the loss of points in the training set which are common to primary and secondary tasks. We emphasize that this resilience is an advantage of the multitask method in comparison with $\Delta$-learning—where no training is feasible without access to secondary C data. Each point on subplot (a) of Fig. 4 compares the average MAE of two multitask models: one with a CA secondary training set, the other with an identical A set but no secondary C data. As before, we are inferring the 3-b energy of water trimer configurations with only one secondary task. The closeness of the points to the reference $y=x$ line shows that there is minimal effect on accuracy when the secondary C data is excluded from the training set. In fact, the median improvement in accuracy from using the CA training model instead of an A training model model is $0.0007$ eV. By comparison, the median improvement in accuracy from combining DFT data from the A set with CCSD(T) training data—that is, switching from a single task approach to a multitask method trained on an A set—is $0.004$ eV.
@@
Subplot (b) of Fig. 4 also compares models trained with and without C secondary data in the case where all of these models also train on secondary Target data. While the AT models perform well, there is a penalty to losing access to secondary C training data when secondary Target data is also available. The median improvement in accuracy when switching from an AT model to a CAT model is $0.002$ eV, a third of the $0.006$ eV median improvement incurred from switching from a single task approach to an AT multitask model with the same CCSD(T) training cost. As we remarked previously, when the model trains on secondary C data, it likely gains information about the difference between the primary and secondary tasks. This information is most useful to the model when it also has access to secondary predictions for the Target molecules.

### v2

@@ III.3 Multiple Secondary Tasks
To make this discussion more tractable, we introduce notation which allows us to compare the sets of molecules used to train each task. Table 2 defines three set-valued functions: $\mathcal{C}$, $\mathcal{A}$, and $\mathcal{T}$. To illustrate this notation further, Figure 2 provides a visual representation of two possible training set structures. Each line in the figure corresponds to a different task, here defined by the level of theory employed during data generation. CCSD(T) informs the primary task, and different DFT functionals correspond to each secondary task (shaded in grey). Black dots represent training data, and their location indicates whether molecules from a given set are available for a particular task. Vertically aligned dots indicate the presence of the same molecular configuration in the training set for multiple tasks. For example, consider the top plot of Figure 2: SCAN and PBE are the secondary tasks, $s_{1}$ and $s_{2}$, and we have:
$\bigl{|}\mathcal{C}(s_{n})\bigr{|}$ $=$ $\ 0,\quad n\in{1,2}$
|$\frac{\bigl{|}\{\mathcal{A}(s_{1})\cap\mathcal{A}(s_{2})\}\bigr{|}}{|A|}$ $=$ $\ 1,$
where ||$\cdot$|| ⋅ | returns the number of elements in a given set, and $A$ is the entire ‘additional’ set. The training data for both secondary tasks only includes molecules from the A set, so this plot represents an A test case. The lower plot represents a CAT data structure where  $\forall\,1\leq n<n^{\prime}\leq 4$ and
|$\frac{\bigl{|}\{\mathcal{C}(s_{n})\cap\mathcal{C}(s_{n^{\prime}})\}\bigr{|}}{|C|}\ =\ \frac{\bigl{|}\{\mathcal{A}(s_{n})\cap\mathcal{A}(s_{n^{\prime}})\}\bigr{|}}{|A|}\ =\ \frac{2}{10}.$
III.4 Training with Target Molecules
We stress that whenever $\mathcal{T}(s_{n})\neq\emptyset$, the secondary training data includes low level predictions of at least some molecules which we will use to test the multitask model. In Fig. 2, the testing molecules—target molecules at the primary fidelity—are represented in green. The black dots in the third column of the lower plot indicate that BLYP predictions for the test molecules are included in the training set.
All lower level predictions for target molecular configurations which are available at training time can be used to obtain a multitask model. If new lower level T predictions become available after this initial training, a retraining is needed to use this data. Note that a $\Delta$-learning model can leverage such additional T data without any retraining. However, each target data point will only contribute to a single prediction of the $\Delta$-learning framework. In contrast, retraining the multitask model buys the benefit of this additional data point for all future predictions. Furthermore, sparse update techniques such as the Sherman-Morrison formula offer efficient alternatives to a full retraining of the multitask model.
@@ IV.1.1 Data Generation
Multitask models can be trained on heterogeneous data sets constructed opportunistically from multiple sources without the need to specify a predefined accuracy hierarchy (apart from singling out the primary task). Our test cases were selected to explore this flexibility by tapping mostly on previously published data sets supplemented with only few additional computations.
For our first test case—inference of 3-b interaction energy—we consider 5986 water trimer configurations and corresponding CCSD(T) calculations. These were randomly selected from 45,332 configurations available from https://github.com/jmbowma/q-AQUA Yu et al. (2022) to obtain training data sizes approximately matching our second test case. To illustrate that even such truncated datasets can be employed to yield reasonable predictive qualities with multitask models, we enrich this data set again by running a small number of cheap density-functional theory simulations: on 2992 of these configurations we employ Psi4 Smith et al. (2020) to compute the 3-b interaction energies at PBEPerdew, Burke, and Ernzerhof (1996) and SCANSun, Ruzsinszky, and Perdew (2015) level. The former was selected for its popularity and the latter because its description of intermediate range dispersion has been found to contribute to accurate prediction of water energy differences Chen et al. (2017); Dasgupta et al. (2021) Following recent work on dispersion corrections, Price, Bryenton, and Johnson (2021) we paired the PBE approximation with a D3 correction using finite damping and the SCAN approximation with a D3 correction using zero damping. Both DFT and CCSD(T) calculations employed a standard counterpoise correction to alleviate basis set superposition error. Huang, Braams, and Bowman (2006)
@@
Quantification of cost in our numerical results refers to the cost of data set generation. The cost of inference is negligible next to quantum chemistry calculations. We base our cost model for each method on the average runtime of Psi4 computations for ten representative systems. For more details about the cost model, see Appendix B. We make the code required to reproduce our results and figures openly available as described in Statement Data Availability Statement. Similarly, the first principles data as well as surrogate prediction data are openly accessible along with the cost of training and applying surrogate models.
@@ IV.1.2 Features
We require features which can distinguish between molecular systems to serve as inputs, $\bm{X}_{i}$, to our Gaussian process models. Kernel functions use features to model the relationships between quantities of interest. As noted in Section II.1, in our experiments we consider Smooth Overlap of Atomic Positions (SOAP) features, which have been successfully used to fit interatomic potentials in a Gaussian process based approach.  Bartók et al. (2010); Bartók, Kondor, and Csányi (2013); Bartók and Csányi (2015) SOAP features are constructed based on the neighborhood of each atom in a system, so challenges arise when two systems contain different numbers of constituent atoms of elements. Multiple strategies have been proposed for constructing “global” features to capture entire systems. De et al. (2016) The simplest approach is to take an average over the SOAP features for each atom in the system. Since this approach is computationally efficient and works well in practice, we apply it in our test cases. We chose SOAP parameters following experimentation and the advice of literature.  Deringer et al. (2021); Musil et al. (2021) More details are available in Appendices C and D.
@@ IV.1.3 Hyperparameter Optimization
Squared exponential kernels are used within all Gaussian process priors in our experiments: Abrahamsen (1997)
$k(\bm{X},\bm{X}^{\prime})\ =\ v\ \exp\left(\frac{-(\bm{X}-\bm{X}^{\prime})^{T}(\bm{X}-\bm{X}^{\prime})}{2\ell}\right)$
These kernels introduce mean, variance ($v$), and length scale ($\ell$) hyperparameters in addition to the multitask correlation ($\rho$). We use the mean of our training data for each task as that task’s prior mean, calculated at training time. To estimate the remaining parameters, we construct an optimization data set containing $150$ randomly selected molecular configurations which we subsequently exclude from our training and testing data sets. We assume here that we have access to predictions of our quantities of interest from each quantum chemistry method for each of these $150$ molecular configurations, but when this assumption does not hold, a conventional GP regression model may be used to fill in missing data.
We estimate correlation parameters, $\rho$, using Pearson’s correlation coefficient between each secondary task optimization data set and the corresponding primary task data. Popular belief in literature holds that correlation parameters require a detailed optimization procedure, Dral, Zubatiuk, and Xue (2023) but we find that these easy to obtain estimates work reasonably well in practice. A more detailed optimization is left for future work. We estimate $\ell\in\mathbb{R}$ by maximizing its log-marginal likelihood. This approach is easily extended to multifidelity and multitask cases by treating $Y_{ij}-\rho_{j}Y_{i1}$ as the $i^{th}$ observation. Forrester and A. Sóbester (2007) For the organic molecules example, we also use a maximum likelihood estimate for $v$ while for the water trimer example, we find better results from using an estimate of $v$ which minimizes mean absolute prediction error. See Appendix A for a comparison of MLE and minimum MAE estimates as well as a more detailed discussion of the resources required for hyperparameter optimization.
IV.1.4 Isolating the Impact of Target Training Data
Consider a molecular configuration, $\bm{X}_{m}$. We are interested in contrasting the accuracy of a multitask model trained using a CA set and a model trained on a CAT set which is identical except for the inclusion of a DFT prediction for $\bm{X}_{m}$. The Target set may contain a large number of molecular configurations, and we cannot make a fair comparison if we simultaneously include DFT predictions for all Target molecules in the training set. To facilitate comparison, whenever $|\mathcal{T}(s_{n})|>0$ for $n=1,\dots,N-1$ we train a new multitask model for each element of $\mathcal{T}(s_{n})$. Each model will train on only one configuration from the $T$ set, and we will test that model by making predictions for the same configuration. We report the mean absolute error (MAE) over all these models and compare this value to the MAE of the predictions of a CA trained model for the same set of targets.
We implement this retraining procedure to investigate the impact of training set design. Retraining is not a prerequisite for leveraging all molecules in $\mathcal{T}(s_{n})$. All lower level data for target molecules which is available at training time can be simultaneously incorporated into one multitask model. See Subsection III.4 for further discussion.
@@ IV.2.1 Multitask Performance
Figure 3: Water Trimer Case. Comparison of the MAE of the multitask method for a range of training set constructions (scatter points) to the MAE of a GP regression model trained only with CCSD(T) data (black line). Predictions by the multitask method are informed by CCSD(T) and SCAN. The legends of the plots indicate the structure of the secondary training set according to the system presented by Fig. 1. The red arrows emphasizes the order of magnitude improvement in data generation cost achieved by CAT multitask models, quantified in Table 3.
We first compare the performance of multitask models with only one secondary task ($N=2$) to the performance of a traditional GP model ($N=1$) . Fig. 3 shows the error incurred by several implementations of each model when predicting the three-body (3-b) energy of water trimer configurations. We used the SCAN functional to generate the secondary task data. Pearson’s correlation coefficient between the primary and secondary task training data is $0.997$, indicating strong positive correlation. The three subplots of Fig. 3 cover six secondary training set configurations (C, CT, A, AT, CA, and CAT), labeled within the legends.
All models were tested on a target set of 320 molecular configurations. For multitask models, we consider seven possible sizes of the core set: $|C|\in\{5,10,20,40,80,160,320\}$. We use the same seven core sets as well as sets with $|C|\in\{640,1280,2560\}$ to train the corresponding single task ($N=1$) models. Performance of this single task reference is marked with a black line on each subplot. Larger $|C|$ is considered in the single task case to demonstrate the cost required for a these models to reach the accuracy of the best performing multitask models. We choose the additional training sets for multitask models so that $|A|=r|C|$ where $r$ is allowed to take on each value in $\{0.5,1,2,3,4,5,6\}$. Whenever core data is included in the secondary training set, we only consider $\mathcal{C}(s_{1})=C$. Thus, we consider a total of $7$ different training set sizes for the C and CT structures and $49$ different training set sizes for the A, AT, CA, and CAT structures. Each point on each subplot of Fig. 3 represents a model with a different training set size. The x axis indicates the cost of each model’s training set; details of our cost model are provided in Appendix B. The y axis gives the average value of mean absolute error obtained from six models trained on a given training set size and structure. For each of these six implementations, available molecular configurations are randomly assigned to the C, A, and T sets without replacement.
For every secondary training set structure except C, the multitask method outperforms the single task reference. Including T data from DFT in the training set leads to better results than corresponding cases without T in every example. In particular, training with secondary T data can enable the multitask method to exhibit an accuracy comparable to a GP model that is an order of magnitude more expensive. Table 3 provides representative examples. The data points corresponding to the last line in the table are marked with arrows on Fig. 3. This analysis of MAE demonstrates that each statistical model can achieve chemical accuracy. For insight into prediction for individual molecules, we report that the best performing multitask models produce predictions with $\geq 95\%$ correlation to the CCSD(T) target data.
Accuracy [meV] Cost [millions of seconds]
MAE Single Task CA CAT
$14$ $2.0$ $0.58$ $0.13$
$7$ $16.0$ $4.5$ $2.0$
Table 3: Water Trimer. Data generation cost to achieve the threshold of accuracy given in the left column.
@@
Figure 4: Water Trimer Case. Effect of including Core and Additional data (defined in Table 1) in secondary training sets. Each indigo point represents a comparison between two multitask models where one model uses a training set which is a subset of the other. In (a), this comparison is between A and CA training configurations, and in (b), we compare AT and CAT configurations. For each pair of models compared, both the primary training data and the secondary data from the A set are identical. The $y=x$ line is plotted in green for reference.
@@
While the C results demonstrate no significant improvement on the single task model, both A and CA results show steady improvement as the size of the A set increases. We can distinguish the impact of increasing $|A|$ from the impact of increasing $|C|$ in Fig. 3 because CCSD(T) is significantly more expensive than DFT. Thus, every jump in $|C|$ leads to a significant increase in cost while jumps in $|A|$ for fixed $|C|$ produce a comparatively slight increase to cost. The positive correlation between $|A|$ and performance whenever $\mathcal{T}(s_{1})=\emptyset$ suggests that for the multitask method to be beneficial, secondary training data must cover molecular space that is not included in the CCSD(T) training data. We mention that it is possible that the behavior of C models would improve if we consider task dependent features or different relationships between task regression functions. We do not investigate these alternatives further here; instead, we focus on the usual setting where secondary DFT data for additional systems is cheap to obtain.
@@
CT and CAT models are the best-performing multitask models, demonstrating very similar MAE per cost to each other. Furthermore, unlike the AT model, the CAT model does not show improvement as the size of the A set increases. It seems that for our setting, the combination of secondary C and T training data is so valuable to the model that any benefit from the A set is dwarfed. We can hypothesize that when the model has access to training data for the Core molecules from both the primary and secondary tasks, it implicitly learns information about the difference between the tasks. Given secondary data on the Target, it can leverage the captured relationship between tasks to make an accurate prediction for the Target on the primary fidelity.
@@
Fig. 4 provides more insight to the relative roles of the training data sets and to the inherent flexibility of the multitask framework. While Fig. 3 demonstrates that multitask models outperform single task models for several training set configurations, it does not directly compare the efficacy of the A and CA training structures. Such a comparison will illuminate the resilience of the multitask model to the loss of points in the training set which are common to primary and secondary tasks. We emphasize that this resilience is an advantage of the multitask method in comparison with $\Delta$-learning—where no training is feasible without access to secondary C data. Each point on subplot (a) of Fig. 4 compares the average MAE of two multitask models: one with a CA secondary training set, the other with an identical A set but no secondary C data. As before, we are inferring the 3-b energy of water trimer configurations with only one secondary task. The closeness of the points to the reference $y=x$ line shows that there is minimal effect on accuracy when the secondary C data is excluded from the training set. In fact, the median improvement in accuracy from using the CA training model instead of an A training model model is $0.0007$ eV. By comparison, the median improvement in accuracy from combining DFT data from the A set with CCSD(T) training data—that is, switching from a single task approach to a multitask method trained on an A set—is $0.004$ eV.
@@
Subplot (b) of Fig. 4 also compares models trained with and without C secondary data in the case where $\mathcal{T}(s_{1})\neq\emptyset$ for all models. While the AT models perform well, there is a penalty to losing access to secondary C training data when secondary Target data is also available. The median improvement in accuracy when switching from an AT model to a CAT model is $0.002$ eV, a third of the $0.006$ eV median improvement incurred from switching from a single task approach to an AT multitask model with the same CCSD(T) training cost. As we remarked previously, when the model trains on secondary C data, it likely gains information about the difference between the primary and secondary tasks. This information is most useful to the model when it also has access to secondary predictions for the Target molecules.